OpenStudy (anonymous):

Ok. So can anyone explain how the proof of 9+10=21 is right or wrong.

OpenStudy (anonymous):

9=3x3 10=5x2 5x3=15 3x2=5 15+6=21

OpenStudy (anonymous):

Or a little easier to see here https://i.ytimg.com/vi/LQpd3bSigiU/hqdefault.jpg

OpenStudy (alexandervonhumboldt2):

Hey, well, you posted this in wrong section.

OpenStudy (alexandervonhumboldt2):

this is broing : 9+10 in real world is 19, in none real world is 21

OpenStudy (anonymous):

oops

OpenStudy (alexandervonhumboldt2):

this is not even interseting

OpenStudy (anonymous):

thanks man

OpenStudy (malcolmmcswain):

9 + 10 = 21 ? Subtract 9 from both sides. 9 - 9 + 10 = 21 - 9 Divide both sides by 1000. 9/1000 - 9/1000 + 10 = 21/1000 - 9/1000 Multiply by x. 9/1000 - 9/1000 + 10x = 21/1000 - 9/1000x Graph.  Now, you can see: that none of this makes sense. :)